OpenStudy (anonymous):

Does each set of numbers form a Pythagorean triple? Explain. 4,5,6 10,24,26 15,20,25

OpenStudy (help!!!!):

square them

OpenStudy (anonymous):

heck no \[4^2+5^2\neq 6^2\]

OpenStudy (help!!!!):

a^2+b^2=c^2

OpenStudy (anonymous):

it is an arithmetic check does \(10^2+24^2=26^2\) ? if so, yes if not, no

OpenStudy (anonymous):

simple check if three numbers are sides of right triangle via Pythagorean theorem. Pick c as biggest number, a,b- smallest and test if a^2+b^2=c^2 ;)